OpenStudy (anonymous):

-8x+4=28 solve for y

OpenStudy (anonymous):

where is the y?

OpenStudy (anonymous):

Ah, sorry. That was suposed to be -8x+4y+28

OpenStudy (anonymous):

-8x +4y=28 28+8x=4y y=(28+8x)/4

OpenStudy (anonymous):

Thanks:)